A radio wave has a frequency of 1x106 Hz. What is it's wavelength?

6A radio wave has a frequency of 1x10


1
Expert's answer
2021-01-12T07:04:48-0500

"f = 1×10^6Hz\\
c = 3 × 10^8 m/s\\
\lambda = ?"


Using,

"c = \lambda f"


"\lambda =\dfrac{c}{f} = \dfrac{3×10^8}{1×10^6} = 3×10^2m" = 300m



Therefore, the wavelength of the radio wave is 300m.